OpenStudy (anonymous):

2x+3y=18 x-y=5 equations of elimination. show workk!

OpenStudy (anonymous):

We can multiple the 2nd equation by 2 and get 2x-2y=10. Then subtract that whole equation from the first 2x+3y = 18 -2x+2y=-10 ______________ 5y = 8 y= 8/5 Now substitute into one of the first equations and solve for x :) Enjoy.

OpenStudy (anonymous):

x=66/5;y=8/5